OpenStudy (dmezzullo):

f f(x) = −4x − 2, find f(−3) 10 −14 12 −10

OpenStudy (anonymous):

plug in -3 for x

OpenStudy (dmezzullo):

@kelly226

OpenStudy (anonymous):

x= 3f/4 - 1/2

OpenStudy (anonymous):

f(x) = −4x − 2 f(-3) = −4(-3) − 2 =?

OpenStudy (anonymous):

f(x) means a function in x. F(x) = -4x-2. f(-3) means replacing value of x with -3 therefore, f(-3)= -4(-3)-2 = 12-2 = 10

OpenStudy (dmezzullo):

tyvm

OpenStudy (anonymous):

np:P

OpenStudy (anonymous):

all you do is plug in -3 wherever there is an x... :)

OpenStudy (anonymous):

all you do is plug in -3 wherever there is an x... :)